# Erasing Self-Supervised Learning Backdoor by Cluster Activation Masking

### v1

@@ Abstract
Researchers have recently found that Self-Supervised Learning (SSL) is vulnerable to backdoor attacks. The attacker can embed hidden SSL backdoors via a few poisoned examples in the training dataset and maliciously manipulate the behavior of downstream models. To defend against SSL backdoor attacks, a feasible route is to detect and remove the poisonous samples in the training set. However, the existing SSL backdoor defense method fails to detect the poisonous samples precisely. In this paper, we propose to erase the SSL backdoor by cluster activation masking and propose a novel PoisonCAM method. After obtaining the threat model trained on the poisoned dataset, our method can precisely detect poisonous samples based on the assumption that masking the backdoor trigger can effectively change the activation of a downstream clustering model. In experiments, our PoisonCAM achieves 96% accuracy for backdoor trigger detection compared to 3% of the state-of-the-art method on poisoned ImageNet-100. Moreover, our proposed PoisonCAM significantly improves the performance of the trained SSL model under backdoor attacks compared to the state-of-the-art method. Our code, poisoned datasets, and trained models will be open once this paper is accepted.
1 Introduction
@@
In this paper, we propose PoisonCAM, which aims at accurately detecting poisonous samples in a poisoned and unlabeled dataset to erase the SSL backdoor. To retrieve the trigger patches injected into the poisoned dataset, we propose a novel Cluster Activation Masking method. Our method is based on the assumption that masking a successful trigger in an image will change the cluster assignment from the target category of the trigger to the true category of the image. Therefore, the trigger patches can be detected by comparing the cluster activation under a few random masks. As shown in Figure 1, our PoisonCAM significantly improves the top 20, 50, and top 100 accuracy of the retrieved trigger patches from 10%, 6%, and 3% of the state-of-the-art PatchSearch to 95%, 96%, and 96% on the poisoned ImageNet-100. Based on the more accurate trigger patches, our trained poison classifier also significantly improves the precision of detecting poisonous samples from 5.4% to 49.3% while achieving even higher recall (see details in Table 2). Finally, after accurately detecting and removing the poisonous samples, the trained SSL model on the clean-up dataset achieves new state-of-the-art results in terms of both the performance and the backdoor defense.
@@
2 Related Work
SSL Backdoor Attacks. The purpose of SSL backdoor attacks is to inject a stealthy backdoor trigger into SSL models by poisoning training data, which can be activated to manipulate the behavior of downstream models at test time [21, 48]. Saha et al. [43] propose backdoor attacks towards SSL models by attaching a trigger patch to images of a target category. At test time, the downstream classifier has high accuracy on clean images but misclassifies images with the trigger as the target category. The authors also propose a distillation-based defender, which requires a substantial amount of trusted data and may be infeasible in real-world scenarios. Li et al. [27] recently propose a similar SSL backdoor attack method. The key difference is that they adopt a frequency domain based spectral trigger to resist the data augmentations. Carlini and Terzis [2] propose backdoor attacks towards CLIP [39], a multimodal contrastive SSL model. Their attack is implemented by injecting triggers into the victim images and tampering with the paired textual captions.
3 Threat Model
@@
4 Proposed Defender Model
@@
4.1 Clustering Model Learning
@@
4.2 Candidate Trigger Detection for Images
Refer to caption (a) 0-1 Interval
Refer to caption (b) Random
Refer to caption (c) Full Coverage
@@ 4.2 Candidate Trigger Detection for Images
Specifically, using the obtained clustering model $C(\cdot)$ in Equation 1, we aim at locating the region of the potential trigger in $x_{i}$. For any image $x_{i}\in{\boldsymbol{X}}$, we randomly initialize $B$ masks $\boldsymbol{M}=\{\boldsymbol{m}_{j}\in\{0,1\}^{H\times W}\}_{j=1}^{B}$. In particular, we set the pixel value to zero if the corresponding mask value is 1 and keep the pixel value unchanged if the corresponding mask value is 0. After randomly selecting a masking window of the size of $[w,w]$, we have empirically tested three masking strategies, as shown in Figure 3. Although it seems that 0-1 Interval Masking and Random Masking can break the pattern of the trigger while protecting the semantics of benign regions, we empirically found in experiments (see Section 5.7) that Full Coverage Masking is the optimal strategy. We apply masks $\{\boldsymbol{m}_{j}\}_{j=1}^{B}$ to obtain $B$ degraded images $\boldsymbol{D}=\{x_{i}^{j}=x_{i}\odot(1-\boldsymbol{m}_{j})\}_{j=1}^{B}$, where $\odot$ denotes the element-wise multiplication. Then we classify the features of $\boldsymbol{D}$ by the clustering model $C(\cdot)$ to get cluster labels $\{\eta_{j}=C(f(x_{i}^{j}))\}_{j=1}^{B}$ and the distances to the assigned cluster centers $\{d_{j}\}_{j=1}^{B}$. Since the size of the backdoor trigger is much smaller than the whole image, most masks cannot change the clustering assignment of $x_{i}^{j}$. Therefore, most samples from $\boldsymbol{D}$ are assigned to the same cluster and we denote this cluster as $\eta^{*}$. To scale cluster distances, we utilize Max-Min Scaling to normalize $d_{j}$ as $d^{n}_{j}=\frac{d_{j}-\min_{k}d_{k}}{\max_{k}d_{k}-\min_{k}d_{k}}\in[0,1]$. Then, we compute the clustering outlier score $a_{j}$ of the mask $\boldsymbol{m}_{j}$ as follows:
@@
4.3 Poisonous Image Search
@@
To find a few highly poisonous images, we can compute the poison scores of all images and simply take the top-$k$ image as the poisonous samples. However, to improve the efficiency and utilize the clustering information of images, we adopt a heuristic search strategy following [49] to iteratively compute poison scores for a part of (not all) images in $\boldsymbol{X}$. Finally, we take the top-$k$ scored images with the highest poison scores to form a poison set $\boldsymbol{X}^{p}$. This poison set will be utilized to train a classifier of poisonous samples.
4.4 Poison Classifier Training
5 Experiments
We include implementation details and more experimental results in Supplementary Material.
5.1 Datasets
Following previous work [49], we adopt ImageNet-100 [50], which contains images belonging to 100 randomly sampled classes from the 1000 classes of ImageNet [41]. The training set has about 127K samples and the validation set has 5K samples. STL-10 [11] contains 500/800 training/validation images for each of 10 classes.
5.2 Attack Setting
Table 1: Defense results. We compared the performance of the attacked SSL models with different defense methods under various attack settings on the validation set. ACC denotes accuracy. ASR denotes attack success rate. Clean Val denotes the clean validation set and Poisoned Val denotes the poisoned validation set.
@@ 5.2 Attack Setting
5.3 Baseline Method and Evaluation Metrics
We adopt the state-of-the-art SSL defense method PatchSearch [49] and the naive method without defense as the baseline methods. We fairly compare our proposed PoisonCAM with baselines by sharing the weights of the threat model. For the evaluation, we train a linear classifier on a trained SSL model by randomly sampling a $1.0\%$ subset from the clean labeled training dataset, following PatchSearch. Since PatchSearch has not released its trained models, we reimplement PatchSearch using the official code [^1] https://github.com/UCDvision/PatchSearch. We will release our poisoned datasets, code, and trained models once this paper is accepted.
We denote the original validation set of ImageNet-100 as the clean validation set, which consists of 50 randomly selected images from each category, and form a poisoned validation set by randomly pasting the attacked trigger on the images of the clean validation set. Similarly, we construct the clean and poisoned validation sets of STL-10. We then assess the results both on the clean validation set and the poisoned validation set by using the following metrics: Accuracy (ACC) and Attack Success Rate (ASR) [55, 59, 28] for the attacked class. Attack Success Rate (ASR) refers to the proportion of non-target classes misclassified by the classifier as the targeted class. All results are averaged over 5 independent runs with different seeds.
5.4 Results and Discussions
The results on poisoned ImageNet-100 and STL-10 under different attack settings are shown in Table 1. Based on the results, we have the following observations:
Table 2: Poison detection results. We compare the average results over target categories for detecting poisonous images in the poisoned datasets. Total Rem. denotes the number of total removed samples.
@@ 5.4 Results and Discussions
5.5 Analysis of Poisonous Image Detection
A key step in PatchSearch and our PoisonCAM is to detect poisonous images in the training set and remove them. To further investigate the effectiveness of our method, we analyze the results of poisonous image detection by PoisonCAM and PatchSearch. In Table 2, we report the number of total removed images (Total Rem.), recall, and precision on three datasets. We have the following observations:
@@ 5.5 Analysis of Poisonous Image Detection
We include the detailed results of every target category in Supplementary Material.
Refer to caption (a) 2-target attack
Refer to caption (b) 3-target attack
Figure 4: Results of the detected top-$k$ candidate triggers by PatchSearch and our PoisonCAM against multi-target attacks.
5.6 Results of Multi-Target Attacks
@@
Refer to caption (a) IoU & CR
Refer to caption (b) Recall & Precision
Refer to caption (a) IoU
Refer to caption (b) CR
Refer to caption (c) F1-score
Refer to caption (d) Average Processing Time
5.7 Ablation Study
To further investigate the effectiveness of the designed masking strategies in Section 4.2, we design different masking strategies and conduct an ablation study on ImageNet-100 (poison rate 0.5%) with the target category “rottweiler”: Recalling that we set the size of masking windows as $[w,w]$, we compare three masking strategies:
@@ 5.7 Ablation Study
5.8 Analysis of Trigger Attention Maps
To further investigate the process of two defenders, PoisonCAM and PatchSearch, we conduct a qualitative analysis of the generated trigger attention maps. Both methods generate trigger attention maps for images and select the hottest regions as the candidate triggers. PatchSearch adopts Grad-CAM [45] to compute the trigger attention, while we propose a Cluster Activation Masking method in Section 4.2 to solve this problem. Therefore, we visualize the cases of trigger attention maps generated by PoisonCAM and PatchSearch on ImageNet-100 (poison rate 0.5%) with the target category “rottweiler” in Figure 6. Our method can always accurately locate the backdoor triggers contained in these images, while PatchSearch focuses on more dispersed and maybe irrelevant regions. These results further demonstrate that our PoisionCAM can significantly improve the detection accuracy of the injected backdoor trigger, which is fundamental for defending against SSL backdoor attacks.
5.9 Hyper-parameter Sensitivity Study
In this section, we analyze the number $B$ of masks in terms of sensitivity with the search number from top-20 to top-100 on ImageNet-100 (poison rate 0.5%) with the target category “rottweiler”. As demonstrated in Figure 7, we have the following observations: (1) While increasing the number $B$ of masks, IoU, CR of the detected candidate triggers and F1-score of the trained poison classifier first increase and then become relatively stable, which shows that our method can effectively locate the triggers with appropriate large $B$. The searching time increases with more masks to calculate. (2) The average processing time consistently increases while the number $B$ of masks increases. Considering the worst situation where all masks are processed serially, the slope of the time curve will tend towards a constant. To conclude, an appropriate value of $B$ should be chosen to balance the performance and time costs. We empirically set $B=256$ in our method.
6 Conclusion
Appendix A Preliminary: Heuristic Search Strategy used in Section 4.3
In Section 4.3, to improve the efficiency and utilize the clustering information of images, we adopt a heuristic search strategy following [49] to iteratively compute poison scores for a part of (not all) images in $\boldsymbol{X}$. To make our paper self-contained and facilitate the comprehension of experimental results, we introduce the algorithm in this section. More details can be found in [49] and our released code.
At every iteration, we first compute the poison scores of $s$ randomly selected samples per cluster and use the sums to represent the poison scores of clusters. Since clusters with lower poison scores are less likely to be poisoned, we then remove a fraction $r$ of the clusters with the least poison scores and will not compute poison scores for images in these clusters. After several iterations, all clusters are removed, and a part of images in $\boldsymbol{X}$ are scored during the iterative procedure. Finally, we take the top-$k$ scored images with the highest poison scores to form a poison set $\boldsymbol{X}^{p}$. This poison set will be utilized to train a classifier of poisonous samples. In this heuristic algorithm, the number of scored images is unfixed.
Appendix B Implementation Details
We utilize PyTorch [37] to implement all experiments on four GeForce RTX 3090 GPUs. We employed the following parameters for PoisonCAM on ImageNet-100 as same as PatchSearch: a cluster count of $l=1000$, the number of the flip test set is set as $1000$, samples per cluster $s=2$, and removal of $r=25\%$ of candidate clusters after each iteration. The search window size was set to $w=60$, using a complete sampling approach. This method typically involves searching through approximately 8,000 images on ImageNet-100. For training the poison classifier, we sample top-20 poisonous images (i.e., $|\boldsymbol{X}^{p}|=20$) and remove the top $10\%$ poisonous samples in $\boldsymbol{X}$ to reduce noise in the poison classification set $\bar{\boldsymbol{X}}$. Besides, we set the number of masks as $B=256$.
Table 3: The attack effectiveness throughout the backdoored SSL process. We evaluated the attack effectiveness using poisoned data to train the self-supervised model with different training epochs. We test the accuracy and false positive rates of the model on clean and poisoned data after fine-tuning on downstream classification tasks using clean datasets. (poison rate 0.5%, target category “laptop” and “goose”). The results show a strong correlation between the effectiveness of the attack and the model performance before 100 epochs.
@@ Appendix B Implementation Details
Table 4: Poison detection results. We compared the results of using both methods for cleaning poisoned data at the data level. Total Rem. denotes the total removed samples.
Attack Target PatchSearch PoisonCAM
[…]
ImageNet-100 (poison rate 0.5%) rottweiler 8449 97.1 7.5 1556 99.5 41.6
tabby cat 11341 99.2 5.7 2314 99.8 28.0
ambulance 13212 31.5 1.6 3509 98.5 18.2
pickup truck 16523 96.8 3.8 921 98.5 69.5
laptop 14944 68.5 3.0 1245 98.2 51.2
goose 14002 99.1 4.6 844 99.1 76.3
pirate ship 18421 53.5 1.6 1924 96.1 27.9
gas mask 4785 99.7 13.5 1457 99.5 44.4
vacuum cleaner 6600 98.9 9.7 806 98.0 79.0
American lobster 19798 99.7 3.3 1132 99.4 57.1
average 12807.5 84.4 5.4 1570.8 98.7 49.3
ImageNet-100 (poison rate 1.0%) rottweiler 20585 99.5 6.3 2083 99.6 62.2
tabby cat 3806 99.8 34.1 5106 99.5 25.3
ambulance 14627 96.4 8.6 1980 97.5 64.0
pickup truck 7918 99.6 16.4 1937 99.4 66.7
laptop 7918 98.2 21.7 1767 99.6 73.3
goose 34793 99.5 3.7 1517 98.9 84.8
pirate ship 10102 99.5 11.0 1925 98.2 57.0
gas mask 6816 99.2 18.7 1961 99.7 66.1
vacuum cleaner 2023 98.8 63.5 1696 99.5 76.3
American lobster 5194 99.8 25.0 1896 98.8 67.8
average 11174.9 99.0 20.9 2186.8 99.1 64.4
STL-10 (poison rate 5.0%) bird 306 98.8 80.7 295 98.8 83.7
car 1755 100.0 14.2 441 98.4 55.8
deer 261 96.8 92.7 260 100.0 96.2
average 744.0 98.5 62.5 332.0 99.1 78.6
Appendix C Additional Experiments
C.1 Detailed Results of Poisonous Image Detection
In this section, we list detailed results for poisonous image detection. As shown in Table 4, our PoisonCAM significantly outperforms PatchSearch for defending every category except target category tabby cat on ImageNet-100 with poison rate $1.0\%$. These results demonstrate that PoisonCAM can enhance the accuracy of detecting poisonous images while reducing the number of erroneously removed benign samples. As a result, PoisonCAM can facilitate more comprehensive training, leading to improved performance of the Self-Supervised Learning (SSL) models.
C.2 More Trigger Attention Maps of Different Target Categories

### v2

@@ Abstract
Self-Supervised Learning (SSL) is an effective paradigm for learning representations from unlabeled data, such as text, images, and videos. However, researchers have recently found that SSL is vulnerable to backdoor attacks. The attacker can embed hidden SSL backdoors via a few poisoned examples in the training dataset and maliciously manipulate the behavior of downstream models. To defend against SSL backdoor attacks, a feasible route is to detect and remove the poisonous samples in the training set. However, the existing SSL backdoor defense method fails to detect the poisonous samples precisely. In this paper, we propose to erase the SSL backdoor by cluster activation masking and propose a novel PoisonCAM method. After obtaining the threat model trained on the poisoned dataset, our method can precisely detect poisonous samples based on the assumption that masking the backdoor trigger can effectively change the activation of a downstream clustering model. In experiments, our PoisonCAM achieves 96% accuracy for backdoor trigger detection compared to 3% of the state-of-the-art method on poisoned ImageNet-100. Moreover, our proposed PoisonCAM significantly improves the performance of the trained SSL model under backdoor attacks compared to the state-of-the-art method. Our code, data, and trained models will be open once this paper is accepted.
Index Terms:
Backdoor attack, self-supervised learning, artificial intelligence security.
I Introduction
@@
In this paper, we propose PoisonCAM, which aims at accurately detecting poisonous samples in a poisoned and unlabeled dataset to erase the SSL backdoor. To retrieve the trigger patches injected into the poisoned dataset, we propose a novel Cluster Activation Masking method. Our method is based on the assumption that masking a successful trigger in an image will change the cluster assignment from the target category of the trigger to the true category of the image. Therefore, the trigger patches can be detected by comparing the cluster activation under a few random masks. As shown in Figure 1, our PoisonCAM significantly improves the top 20, 50, and top 100 accuracy of the retrieved trigger patches from 10%, 6%, and 3% of the state-of-the-art PatchSearch to 95%, 96%, and 96% on the poisoned ImageNet-100. Based on the more accurate trigger patches, our trained poison classifier also significantly improves the precision of detecting poisonous samples from 5.4% to 49.3% while achieving even higher recall (see details in Table II). Finally, after accurately detecting and removing the poisonous samples, the trained SSL model on the clean-up dataset achieves new state-of-the-art results in terms of both the performance and the backdoor defense.
@@
II Related Work
SSL Backdoor Attacks. The purpose of SSL backdoor attacks is to inject a stealthy backdoor trigger into SSL models by poisoning training data, which can be activated to manipulate the behavior of downstream models at test time [33, 34, 35]. Saha et al. [12] propose backdoor attacks towards SSL models by attaching a trigger patch to images of a target category. At test time, the downstream classifier has high accuracy on clean images but misclassifies images with the trigger as the target category. The authors also propose a distillation-based defender, which requires a substantial amount of trusted data and may be infeasible in real-world scenarios. Li et al. [36] recently propose a similar SSL backdoor attack method. The key difference is that they adopt a frequency domain based spectral trigger to resist the data augmentations. Carlini and Terzis [11] propose backdoor attacks towards CLIP [37], a multimodal contrastive SSL model. Their attack is implemented by injecting triggers into the victim images and tampering with the paired textual captions.
@@ II Related Work
III Threat Model
@@
IV Proposed Defender Model
@@
IV-A Clustering Model Learning
@@
IV-B Candidate Trigger Detection for Images
(a) 0-1 Interval
(b) Random
(c) Full Coverage
@@ IV-B Candidate Trigger Detection for Images
Specifically, using the obtained clustering model $C(\cdot)$ in Equation 1, we aim at locating the region of the potential trigger in $x_{i}$. For any image $x_{i}\in{\boldsymbol{X}}$, we randomly initialize $B$ masks $\boldsymbol{M}=\{\boldsymbol{m}_{j}\in\{0,1\}^{H\times W}\}_{j=1}^{B}$. In particular, we set the pixel value to zero if the corresponding mask value is 1 and keep the pixel value unchanged if the corresponding mask value is 0. After randomly selecting a masking window of the size of $[w,w]$, we have empirically tested three masking strategies, as shown in Figure 3. Although it seems that 0-1 Interval Masking and Random Masking can break the pattern of the trigger while protecting the semantics of benign regions, we empirically found in experiments (see Section VI-D) that Full Coverage Masking is the optimal strategy. We apply masks $\{\boldsymbol{m}_{j}\}_{j=1}^{B}$ to obtain $B$ degraded images $\boldsymbol{D}=\{x_{i}^{j}=x_{i}\odot(1-\boldsymbol{m}_{j})\}_{j=1}^{B}$, where $\odot$ denotes the element-wise multiplication. Then we classify the features of $\boldsymbol{D}$ by the clustering model $C(\cdot)$ to get cluster labels $\{\eta_{j}=C(f(x_{i}^{j}))\}_{j=1}^{B}$ and the distances to the assigned cluster centers $\{d_{j}\}_{j=1}^{B}$. Since the size of the backdoor trigger is much smaller than the whole image, most masks cannot change the clustering assignment of $x_{i}^{j}$. Therefore, most samples from $\boldsymbol{D}$ are assigned to the same cluster and we denote this cluster as $\eta^{*}$. To scale cluster distances, we utilize Max-Min Scaling to normalize $d_{j}$ as $d^{n}_{j}=\frac{d_{j}-\min_{k}d_{k}}{\max_{k}d_{k}-\min_{k}d_{k}}\in[0,1]$. Then, we compute the clustering outlier score $a_{j}$ of the mask $\boldsymbol{m}_{j}$ as follows:
@@
IV-C Poisonous Image Search
@@
To find a few highly poisonous images, we can compute the poison scores of all images and simply take the top-$k$ image as the poisonous samples. However, to improve the efficiency and utilize the clustering information of images, we adopt a heuristic search strategy following [14] to iteratively compute poison scores for a part of (not all) images in $\boldsymbol{X}$. At every iteration, we first compute the poison scores of $s$ randomly selected samples per cluster and use the sums to represent the poison scores of clusters. Since clusters with lower poison scores are less likely to be poisoned, we then remove a fraction $r$ of the clusters with the least poison scores and will not compute poison scores for images in these clusters. After several iterations, all clusters are removed, and a part of images in $\boldsymbol{X}$ are scored during the iterative procedure. Finally, we take the top-$k$ scored images with the highest poison scores to form a poison set $\boldsymbol{X}^{p}$. This poison set will be utilized to train a classifier of poisonous samples. In this heuristic algorithm, the number of scored images is unfixed. This poison set will be utilized to train a classifier of poisonous samples.
TABLE I: Defense results. We compared the performance of the attacked SSL models with different defense methods under various attack settings on the validation set. ACC denotes accuracy. ASR denotes attack success rate. Clean Val denotes the clean validation set and Poisoned Val denotes the poisoned validation set.
@@
IV-D Poison Classifier Training
V Experiments
Focusing on the self-supervised learning (SSL) backdoor defense task, we conduct extensive experiments on two widely-adopted benchmark datasets.
V-A Datasets
Following previous work [14], we adopt ImageNet-100 [49], which contains images belonging to 100 randomly sampled classes from the 1000 classes of ImageNet [50]. The training set has about 127K samples and the validation set has 5K samples. We also adopt STL-10 [51] that contains 500/800 training/validation images for each of the 10 classes.
V-B Attack Setting
V-C Baseline Method and Evaluation Metrics
We adopt the state-of-the-art SSL defense method PatchSearch [14] and the naive method without defense as the baseline methods. We fairly compare our proposed PoisonCAM with baselines by sharing the weights of the threat model. For the evaluation, we train a linear classifier on a trained SSL model by randomly sampling a $1.0\%$ subset from the clean labeled training dataset, following PatchSearch. Since PatchSearch has not released its trained models, we reimplement PatchSearch using the official code.  [^1] https://github.com/UCDvision/PatchSearch. We will release our poisoned datasets, code, and trained models once this paper is accepted.
We denote the original validation set of ImageNet-100 as the clean validation set, which consists of 50 randomly selected images from each category, and form a poisoned validation set by randomly pasting the attacked trigger on the images of the clean validation set. Similarly, we construct the clean and poisoned validation sets of STL-10. We then assess the results on both the clean validation set and the poisoned validation set by using the following metrics: Accuracy (ACC) and Attack Success Rate (ASR) [55, 56, 57] for the attacked class. Attack Success Rate (ASR) refers to the proportion of non-target classes misclassified as the targeted class. All results are averaged over 5 independent runs with different seeds.
TABLE II: Poison detection results. We compare the average results over target categories for detecting poisonous images in the poisoned datasets. Total Rem. denotes the number of total removed samples.
@@ V-C Baseline Method and Evaluation Metrics
VI Implementation Details
We utilize PyTorch [58] to implement all experiments on four GeForce RTX 3090 GPUs. We employed the following parameters for PoisonCAM on ImageNet-100 as same as PatchSearch: the cluster count of $l=1000$, the number of the flip test set is set as $1000$, samples per cluster $s=2$, and removal of $r=25\%$ of candidate clusters after each iteration. The search window size was set to $w=60$, using a complete sampling approach. This method typically involves searching through approximately 8,000 images on ImageNet-100. For training the poison classifier, we sample top-20 poisonous images (i.e., $|\boldsymbol{X}^{p}|=20$) and remove the top $10\%$ poisonous samples in $\boldsymbol{X}$ to reduce noise in the poison classification set $\bar{\boldsymbol{X}}$. Besides, we set the number of masks as $B=256$.
Given the height $H$ and width $W$ of the images, the $(x,y)$ positions of our sampled masks are in cyclic groups. For a pixel $(x,y)$ in the sampled mask, its real position is $(x\ {\rm mod}\ H,y\ {\rm mod}\ W)$, where the originally sampled positions $x$ and $y$ can exceed $H$ and $W$.
VI-A Results and Discussions
The results on poisoned ImageNet-100 and STL-10 under different attack settings are shown in Table I. Based on the results, we have the following observations:
(a) 2-target attack
(b) 3-target attack
Figure 4: Results of the detected top-$k$ candidate triggers by different methods against multi-target attacks.
(a) IoU & CR
(b) Recall & Precision
(a) ACC $\uparrow$
(b) ASR $\downarrow$
Figure 10: Defense results with i-CutMix augmentation on ImageNet-100 (poison rates 0.5% and 1.0%, target category “rottweiler”).
VI-B Analysis of Poisonous Image Detection
A key step in PatchSearch and our PoisonCAM is to detect poisonous images in the training set and remove them. To further investigate the effectiveness of our method, we analyze the results of poisonous image detection by PoisonCAM and PatchSearch. In Table II, we report the number of total removed images (Total Rem.), recall, and precision on three datasets. We have the following observations:
@@ VI-B Analysis of Poisonous Image Detection
We include the detailed results of every target category in Supplementary Materials.
VI-C Results of Multi-Target Attacks
@@
VI-D Ablation Study
To further investigate the effectiveness of the designed masking strategies in Section IV-B, we design different masking strategies and conduct an ablation study on ImageNet-100 (poison rate 0.5%) with the target category “rottweiler”: Recalling that we set the size of masking windows as $[w,w]$, we compare three masking strategies:
@@ VI-D Ablation Study
(a) IoU
(b) CR
(c) F1-score
(d) Average Processing Time
TABLE III: The attack effectiveness throughout the backdoored SSL process.
@@
VI-E Analysis of Trigger Attention Maps